a person is running at 10km/h. in 5 min he covers a distance of??
500m
600m
833m
1200m
work shown

Found 2 solutions by jorel1380, ewatrrr:
Answer by jorel1380(3719)

10km/hr x 1/12hr=10/12 km=833 m..

Answer by ewatrrr(24785)

 
Hi,
1km = 1000m
a person is running at 10km/hr or 10,000m/60min = 166.67m/min
166.67m/min *5min = 833.33m